# Towards Relaxed Multimodal Inputs for Gait-based Parkinson’s Disease Assessment

### v1

@@ I Introduction
Parkinson’s disease (PD) affects approximately 1–2% of the global population over the age of 65 and is second only to Alzheimer’s disease among neurodegenerative disorders [1]. PD symptoms predominantly include tremor, speech impairment, and gait disturbance, arising from motor neuron degeneration [2]. Symptom severity increases with disease progression; therefore, timely intervention is crucial to prevent deterioration and preserve quality of life [3, 4]. At present, PD assessment is mainly conducted by physicians using well-established clinical scales such as the Movement Disorder Society–Unified Parkinson’s Disease Rating Scale (MDS-UPDRS) and the Hoehn and Yahr (H&Y) scale [2]. However, in the absence of a definitive biomarker, questionnaire-based assessment is time- and resource-intensive, subjective, and insufficiently sensitive to minor fluctuations [5]. There is thus a substantial need to streamline PD detection and progression assessment. With the growing availability of sensing technologies, researchers can now collect high-precision motion data from PD patients, providing rich, objective information that can complement traditional in-clinic evaluation.
@@
Many types of sensor data have been explored for PD assessment, including speech [6, 7], movement [8, 9, 10], neuroimaging [11, 12], and others. Among these modalities, gait-based movement data offer several advantages for diagnosis and severity monitoring [13]. First, gait disturbance is a prominent motor symptom even in the early stages of PD [14]. Second, quantitative gait analysis—explicitly endorsed in clinical guidelines (e.g., Timed Up and Go)—provides objective and reproducible metrics. Moreover, portable devices such as pressure sensors [15], depth sensors [16, 17], cameras [18, 19], their combinations [20, 21], and others[22] can capture multi-perspective, high-precision gait data, enabling a cost-effective yet multidimensional view of PD pathophysiology. Accordingly, we focus on multimodal gait data for PD assessment.
On the other hand, artificial intelligence (AI) algorithms, in particular, have been extensively applied in the existing literature to recognize disease-specific patterns stored in these gait data to aid in the evaluation of PD [23]. The current literature can be categorized into single- and multi-modality ones, where single-modality approaches rely on a single data source (e.g., video[24], skeleton[25], or sensor), while multi-modality approaches integrate multiple data streams (e.g., combining camera and inertial sensors[17]) for a more comprehensive assessment. While single-modality methods have shown competitive performance and currently dominate the field, an increasing body of work has demonstrated the advantages of multi-modality data in PD analytics [23, 5], contributing to the growing popularity of multimodal solutions.
However, existing multimodal solutions rely on a strict data collection process, requiring synchronized inputs: i.e. strictly time-aligned signals from full modalities during both training and inference (see Table I). This requirement presents several key challenges, including technical and personnel-related barriers. Technically, capturing synchronized multimodal gait data (e.g., vertical ground reaction force (vGRF) and skeleton data) often demands specialized equipment and complex, non-standardized experimental setups. Personnel-related issues include data labeling and calibration for synchronizing input data[26]. Moreover, collecting all modalities during inference is often impractical due to privacy concerns and device limitations[27].
@@
In addition, during our experiments with some multimodal fusion models we observed a recurring failure mode which causes the fused model heavily leans on one subset of modalities and if this subset of modalities are removed during inference, the model’s performance significantly degraded. This issue has serious effect on our scenario as missing modality is very common in real-clinic. In the synchronous setting of Table II, ablations make this asymmetry explicit: on WearGait, Early Fusion falls from 82.44% (full) to 50.72% when IMU is removed, yet remains 60.02%/61.72% after removing Insole/Walkway; on FOG, Early/Late/Shared drop to 36.42/36.71/36.14% without Sensor (near 33% chance for three classes), but stay at 40.09/47.64/43.56% without Skeleton. To the best of our knowledge, this phenomenon has not been discussed in gait-based PD assessment. Nevertheless, a subsequent review showed it is not unique to our field and has been discussed in multimodal computer vision field as “modality collapse”, where fusion models rely on a subset of modalities or suppress one branch during optimization [28, 29].
To address above mentioned challenges encountered frequently in real-clinic scenarios, we propose a modality-relaxed multimodal framework that allows asynchronous and optional modality inputs during both training and inference. As a first step, we design a new architecture (see Fig. 1) consisting of modality-specific encoders, a shared feature extractor, and modality-specific prediction heads. This design enables interaction between modalities while preserving modality-specific features. Next, we employ a multi-objective optimization (MOO) algorithm that balances convergence and conflict avoidance to not only facilitate the learning of shared representations across modalities and subjects, but also mitigating the modality collapse issue. We conduct extensive experiments under both synchronous and asynchronous conditions. We showed through experiments that our model also improves single modality accuracy compares to their single-modal baselines. In addition, we introduce a class-rebalancing training scheme to mitigate class imbalance within each modality. In a nutshell, our contributions can be summarized as three-fold:
@@ II Related Work
Our work primarily lies at the intersection of PD assessment and MOO. In the following, we introduce key classical and recent developments in both domains, and then explicitly highlight the connections and differences between our approach and existing research.
@@ II-A Gait-based Single-Modality Solutions
Advancements in sensor devices and AI have enabled the shift in PD assessment from in-clinic evaluations with specialist supervision to AI-enhanced, automated, or remote monitoring approaches [30, 5, 31, 32, 33]. A common direction is to analyze gait data during locomotion, which includes kinematic data (e.g., RGB-D video [16, 17, 18, 19, 20, 21], IMU [34, 35]) and kinetic data (e.g., vGRF [36, 37]). Early ML-based approaches use hand-crafted features (e.g., stride, speed, cadence) with algorithms such as SVM [38], k-NN [39], or ensembles [40, 13], but require extensive manual feature engineering. To reduce this reliance, DL models are introduced to learn representations directly from raw data [41], ranging from MLPs [42] to CNNs [43], LSTMs [44], and GNNs [45], each suited for spatial or temporal signals [46, 47, 48, 13]. More recent methods adopt transformers[49] and hybrid models[50, 37] to capture complex spatiotemporal patterns. Overall, DL has been shown to consistently outperform ML in many PD tasks [51].
@@ II-B Gait-based Multi-Modality Solutions
In addition to single-modality studies, multimodal approaches have demonstrated superior capacity in capturing holistic disease patterns by integrating diverse data sources [5]. However, effectively utilizing cross-modal information remains challenging. Early works either relied on ML with hand-crafted features [52] or combined modality-specific architectures, such as 3D CNNs and LSTMs [53, 54], to extract spatial and temporal features, respectively. While these methods capture partial spatiotemporal or multimodal characteristics, they struggle with cross-modal correlation. More recent efforts can be categorized into three strategies (see Table I): early/late fusion, cross-attention, and shared latent representation. For instance, [17] proposed a late fusion method that uses separate spatial encoders followed by Correlative Memory Neural Networks to learn joint temporal embeddings. The work of [55] introduced a graph-based shared latent fusion by modeling each modality as graph vertices to learn inter-modality relationships. Another work [56] adopted cross-attention modules, applying bi-directional co-attention across silhouette and skeleton streams to extract complementary information.
@@ III Principal Design
In this section, we present a detailed overview of our proposed framework (see Fig. 1), architectural design, MOO learning paradigm, and class re-balancing strategy.
@@ III-A Problem Setup
We consider multimodal gait-based PD assessment with input modalities per subject. Let the dataset be
where denotes the sequence (or window) from modality , is the clinical label. A model consumes any available subset of modalities and predicts the class:
@@ III-B Overall Framework
For each modality , a modality–specific encoder first extracts features and maps to a fixed-width sequence
Then, each is passed independently through a shared encoder (parameters shared across modalities):
The shared representations are then fed to modality–specific heads to produce logits
and per–modality losses are computed. Then, the private parameters are updated with their respective gradients, while a designed MOO algorithm combines the per–modality gradients on the shared encoder to update the shared-backbone ; During the inference phase, optional modalities can be fed into the trained model to obtain their respective outputs.
@@ III-D MOO-based Multimodal Learning
To further enable the asynchronous and optional modality input, we develop a MOO learning paradigm to facilitate the shared feature extraction on shared backbone network () across modalities.
@@
where is the update vector for , i.e., , and is the learning rate. This formulation first seeks the least progress one, and then optimize to improve it. Note that Eqn. 6 can be further transformed as follow by applying first-order Taylor approximation:
where is . Note that , where is a simplex. Then we impose a norm constraint on in Eqn. 7, i.e., . is the average gradient of all modalities, which serves as a naïve optimization. is a hyper-parameter measures the allowed deviation from . This constraint ensures the convergence of MOO algorithm. Therefore, we formulate the dual objective according to Lagrangian equation and Slater’s condition as follow:
where . By fixing and , we can obtain the optimal . Inserting the optimal into Eqn. 8, we have:
@@
Once the optimal are derived according to Eqn. 10, we have , and is updated as follow:
@@
Remarks: Only shared backbone’s parameters are updated via Eqn. 11, other modules are updated by their corresponding modality-specific gradients (see Algorithm 1).
@@ III-E Margin-based Rebalancing Strategy
Another issue lies in the prevalent class imbalance in PD multimodal datasets. As depicted in Fig. 3, both FOG and FBG exhibit a long-tailed distribution across categories. To mitigate the adverse effects of long‐tailed class distributions inherent in PD multimodal gait classification tasks, we integrate class‐adaptive margins and stochastic smoothing into the standard cross‐entropy framework for both modalities.
@@ IV Implementation
IV-A Training Strategy
@@
Crucially, the asynchronous setting is ill-posed for the baseline models as mixing subjects introduces cross-label contamination. We report their asynchronous results only for completeness—to highlight this limitation and to compare against our design, which is expressly built for asynchronous inputs. As shown in Table III, our approach outperforms all baselines by a large margin.
IV-B Validation Strategy
@@
IV-C Sampling Strategy
To mitigate imbalances, we use two complementary strategies. (1) Modality-balanced training (asynchronous only) in asynchronous training, the sample size of each modality varies, so we perform modality-level oversampling. However, this does not guarantee a class-balanced training set within each modality. To address this, we propose a margin-based class rebalancing strategy, which helps ensure class-balanced training. This stabilizes gradient contributions and enhances joint representation learning alongside the MOO objective. (2) Balanced evaluation (both) for fair validation, we oversample the evaluation set to equalize class counts and segment counts within and across modalities.
IV-D Baselines
We evaluate on both single modality and multimodal fusion baselines. All models inherit similar design of encoders, backbone, and training hyper-parameters as our pipeline to isolates fusion effects from all other factors under both synchronous and asynchronous settings.
Single–modality baselines use the exact architecture from our main pipeline, but only one encoder is active at a time. These serve as upper bounds and isolate modality–specific signal.
@@ IV-D Baselines
Conventional methods serve as naive fusion baselines for our approach. (1) Early/Late Fusion: Early Fusion concatenates low-level features from encoders and feeds the joint sequence to the shared backbone; Late Fusion processes each modality independently through the backbone, then concatenates the resulting high-level vectors for prediction. (2) Shared-Latent Fusion: each modality is linearly projected into a common latent width and fused by element-wise addition before the backbone. (3) Cross-Attention Fusion: a lightweight, symmetric cross-attention block lets each modality attend to the other; the fused sequence is then passed to the backbone.
@@
Advanced variants are implemented to benchmark our method against newer designs. The implementation details for below variants are included in the Appendix. (1) FOCAL [69]: factorizes each modality into shared and private latents, enforces orthogonality between them, and applies a contrastive objective to align shared parts across modalities, preserving complementary cues while reducing single-stream shortcuts. (2) TACA [70]: compresses sequences into a small set of learned tokens per modality and performs cross-attention among tokens (plus a time-shared module) for efficient, salient fusion that tolerates time-misaligned streams. (3) DEEPAV [71]: uses lightweight per-modality streams followed by late cross-modal interaction and an agreement/consensus head, emphasizing decision-level fusion and robust fallback when one modality is weak or missing.
IV-E Training and Hyperparameters
Hyperparameter tuning is performed per dataset and model configuration, optimizing for classification accuracy. All models are trained using the Adam optimizer for a maximum of 100 epochs, with early stopping applied after ten epochs without validation performance improvement. The learning rate is fixed throughout with patience set to 50 epochs. Experiments are conducted using an NVIDIA Tesla V100 GPU.
@@ IV-E Training and Hyperparameters
Dataset Approach Modality 1 Modality 2 Modality 3 Average Acc FOG Dataset Skeleton IMU — Skeleton Only - - - - Sensor Only - - - - Early Fusion - Late Fusion - Shared Latent - Cross Attention - FOCAL [69] - TACA [70] - DEEPAV [71] - TRIP - FBG Dataset Skeleton vGRF — Skeleton Only - - - - Sensor Only - - - - Early Fusion - Late Fusion - Shared Latent - Cross Attention - FOCAL [69] - TACA [70] - DEEPAV [71] - TRIP - WearGait Dataset Walkway Insole IMU Walkway Only - - - - Insole Only - - - - IMU Only - - - - Early Fusion Late Fusion Shared Latent Cross Attention FOCAL [69] TACA [70] DEEPAV [71] TRIP
Dataset Sample Count Class Count Modalities FOG [72] 1856 3 (H&Y Scale) IMU Skeleton FBG [73] 4462 3 (MDS-UPDRSIII) vGRF Skeleton WearGait [74] 5219 2 (PD or Healthy) Walkway Insole, IMU
@@
Dataset Approach Accuracy (%) FOG Dataset Skeleton Only – IMU Only – Early Fusion Late Fusion Shared Latent Cross Attention FOCAL [69] TACA [70] DEEPAV [71] TRIP WearGait Dataset Walkway Only – Insole Only – IMU Only – Early Fusion Late Fusion Shared Latent Cross Attention FOCAL [69] TACA [70] DEEPAV [71] TRIP

### v2

@@ I Introduction
Parkinson’s disease (PD) affects approximately 1–2% of the global population over the age of 65 and is second only to Alzheimer’s disease among neurodegenerative disorders [1]. PD symptoms predominantly include tremor, speech impairment, and gait disturbances, arising from motor neuron degeneration. Symptom severity increases with disease progression; therefore, timely intervention is crucial to prevent deterioration and preserve quality of life. At present, PD assessment is mainly conducted by physicians using well-established clinical scales such as the Movement Disorder Society–Unified Parkinson’s Disease Rating Scale (MDS-UPDRS) and the Hoehn and Yahr (H&Y) scale [2]. However, in the absence of a definitive biomarker, the questionnaire-based assessment is time- and resource-intensive, subjective, and insufficiently sensitive to minor fluctuations [3]. Thus, there is a substantial need to streamline PD detection and progression assessment. With the growing availability of sensing technologies, researchers can now collect high-precision motion data from PD patients, providing rich, objective information that can complement traditional in-clinic evaluation.
@@
Many types of sensor data have been explored for the assessment of PD, including speech [4], movement [5], neuroimaging [6], and others. Among these modalities, gait-based movement data offer several advantages for diagnosis and severity monitoring [14]. First, gait disturbance is a prominent motor symptom even in the early stages of PD [8]. Second, quantitative gait analysis—explicitly endorsed in clinical guidelines (e.g., Timed Up and Go)—provides objective and reproducible metrics. Moreover, portable devices such as pressure sensors [9], depth sensors [10], cameras [11], their combinations [12], et al [13] can capture multi-perspective high-precision gait data, enabling a cost-effective yet multidimensional view of the pathophysiology of PD.
On the other hand, artificial intelligence (AI) algorithms, in particular, have been extensively applied in the existing literature to recognize disease-specific patterns stored in these gait data to aid in the evaluation of PD [14]. The current literature can be categorized into single- and multi-modality ones, where single-modality approaches rely on a single data source, while multimodal approaches integrate multiple data streams (e.g., combining camera and inertial sensors[10]) for a more comprehensive assessment. Although single-modality methods have shown competitive performance and currently dominate the field, an increasing body of work has demonstrated the advantages of multi-modality data in PD analytics [3], contributing to the growing popularity of multimodal solutions. Consequently, in this work we focus on multimodal gait data for PD assessment.
However, existing multimodal solutions rely on a strict data collection process, requiring synchronized inputs (i.e., strictly time-aligned signals) from full modalities during both training and inference (see Table I). This requirement presents several key challenges, including technical and personnel-related barriers. Technically, capturing synchronized multimodal gait data (e.g., vertical ground reaction force (vGRF) and skeleton data) often demands specialized equipment and complex, non-standardized experimental setups. Personnel-related issues include data labeling and calibration for synchronizing input data[15]. Moreover, collecting all modalities during inference is often impractical due to privacy concerns and device limitations[16]. Therefore, an approach that handles missing modality and asynchronous streams can greatly alleviate these constraints.
@@
In addition, during our experiments with some multimodal fusion models we observed a recurring failure mode which causes the fused model heavily leans on one subset of modalities and if this subset of modalities are removed during inference, the model’s performance significantly degraded (see Table. II). This issue has serious effect on our scenario as missing modality is very common in real-clinic. In the synchronous setting, ablations make this asymmetry explicit: on WearGait, Early Fusion falls from 82.44% (full) to 50.72% when IMU is removed, yet remains 60.02%/61.72% after removing Insole/Walkway; on FOG, Early/Late/Shared drop to 36.42/36.71/36.14% without Sensor (near 33% chance for three classes), but stay at 40.09/47.64/43.56% without Skeleton. To the best of our knowledge, this phenomenon has not been discussed in gait-based PD assessment. Nevertheless, a subsequent review showed it is not unique to our field and has been discussed in multimodal computer vision field as “modality collapse”, where fusion models rely on a subset of modalities or suppress one branch during optimization [17].
To address above mentioned challenges encountered frequently in real-clinic scenarios, we propose a modality-relaxed multimodal framework that allows asynchronous and optional modality inputs during both training and inference. As a first step, we design a new architecture (see Fig. 1) which enables interaction between modalities while preserving modality-specific features. Next, we employ a multi-objective optimization (MOO) algorithm that balances convergence and conflict avoidance to not only facilitate the learning of shared representations across modalities and subjects, but also mitigating the modality collapse issue. In addition, we introduce a class-rebalanced training scheme to mitigate class imbalance within each modality. Extensive experiments are conducted under both synchronous and asynchronous conditions. The experiment results show that our model not only achieves good accuracy under asynchronous condition, but also improves single modality accuracy compares to their single-modal baselines. In a nutshell, our contributions can be summarized as three-fold:
@@ II Related Work
Our work primarily lies at the intersection of PD assessment and MOO. In the following, we introduce key classical and recent developments in both domains and then explicitly highlight the connections and differences between our approach and existing research.
@@ II-A Gait-Based Single-Modality Solutions
Advancements in sensor devices and AI have enabled the shift in PD assessment from in-clinic evaluations with specialist supervision to AI-enhanced, automated, or remote monitoring approaches [18, 19, 20]. A common direction is to analyze gait data during locomotion, which includes kinematic data (e.g., RGB-D video [10], IMU [21]) and kinetic data (e.g., vGRF [22]). Early ML-based approaches use hand-crafted features (e.g., stride, speed, and cadence) with algorithms such as SVM [23], k-NN [24], or ensembles [7], but require extensive manual feature engineering. To reduce this reliance, DL models are introduced to learn representations directly from raw data [25], ranging from MLPs [26] to CNNs [27], LSTMs [28], and GNNs [29], each suited for spatial or temporal signals. More recent methods adopt transformers[30] and hybrid models[22] to capture complex spatiotemporal patterns. Overall, DL has been shown to consistently outperform ML in many PD tasks [31].
@@ II-B Gait-Based Multi-Modality Solutions
In addition to single-modality studies, multimodal approaches have demonstrated superior capacity in capturing holistic disease patterns by integrating diverse data sources [3]. However, effectively utilizing cross-modal information remains challenging. Early works either relied on ML with hand-crafted features [32] or combined modality-specific architectures (e.g., 3D CNNs and LSTMs [33, 34]), to extract spatial and temporal features separately. While these methods capture partial spatiotemporal or multimodal characteristics, they struggle with cross-modal correlation. More recent efforts can be categorized into three strategies (see Table I): early/late fusion, cross-attention, and shared latent representation. For instance, [10] proposed a late fusion method that uses separate spatial encoders followed by Correlative Memory Neural Networks to learn joint temporal embeddings. The work of [35] introduced a graph-based shared latent fusion by modeling each modality as graph vertices to learn inter-modality relationships. Another work [36] adopted cross-attention modules, applying bi-directional co-attention across silhouette and skeleton streams to extract complementary information.
@@ III Principal Design
In this section, we present a detailed overview of our proposed framework (see Fig. 1), architectural design, MOO learning paradigm, and class rebalancing strategy.
@@ III-A Problem Setup
We consider multimodal gait-based PD assessment with input modalities per subject. Let
where denotes the sequence window from modality and is the clinical label. At training and inference, a subset of modalities may be available due to asynchrony or dropout. A model consumes any available subset and predicts
@@ III-B Overall Framework
For each modality , a modality-specific encoder maps the raw sequence to a fixed-width feature sequence
Each is then processed by a shared encoder to produce shared representations
A modality-specific head produces logits
and a per-modality supervised loss is computed for every . Private parameters are updated using their respective gradients from . Gradients on the shared encoder induced by are combined by a MOO step to yield a single update for . At inference, any subset of modalities can be provided.
@@ III-D MOO-Based Multimodal Learning
To further enable the asynchronous and optional modality input, we develop a MOO learning paradigm to facilitate the shared feature extraction on shared backbone () across modalities.
@@
where is the update vector for , i.e., , and is the learning rate. This formulation first seeks the least progress one, and then optimize to improve it. Note that Eqn. 1 can be further transformed as follow by applying first-order Taylor approximation:
where is . Note that , where is a simplex. Then we impose a norm constraint on in Eqn. 2, i.e., . is the average gradient of all modalities, which serves as a naïve optimization. is a hyper-parameter measures the allowed deviation from . This constraint ensures the convergence of MOO algorithm. Therefore, we formulate the dual objective according to Lagrangian equation and Slater’s condition as follow:
where . By fixing and , we can obtain the optimal . Inserting the optimal into Eqn. 3, we have:
@@
Once the optimal are derived according to Eqn. 5, we have , and is updated as follow:
@@
Remarks: Only shared backbone’s parameters are updated via Eqn. 6, other modules are updated by their corresponding modality-specific gradients (see Algorithm 1).
@@ III-E Margin-Based Rebalancing Strategy
Another issue lies in the prevalent class imbalance in PD multimodal datasets. As depicted in Fig. 3, both FOG and FBG exhibit a long-tailed distribution across categories. To mitigate the adverse effects of long‐tailed class distribution inherent in PD multimodal gait classification tasks, we integrate class‐adaptive margins and stochastic smoothing into the standard cross‐entropy framework for both modalities.
@@ IV Implementation
IV-A Setup
Sampling Strategy
To mitigate imbalances, we use two complementary strategies. (1) Modality-Balanced Training (asynchronous only): in asynchronous training, the sample size of each modality varies, so we perform modality-level oversampling. However, this does not guarantee a class-balanced training set within each modality. To address this, we propose a margin-based class rebalancing strategy, which helps ensure class-balanced training. This stabilizes gradient contributions and enhances joint representation learning alongside the MOO objective. (2) Class- and Modality-Balanced Evaluation (synchronous and asynchronous): for fair validation, we oversample the evaluation set to equalize class counts and segment counts within and across modalities.
Training and Hyperparameters
Hyperparameter tuning is performed per dataset and model configuration, optimizing for classification accuracy. All models are trained using the Adam optimizer for a maximum of 100 epochs and the learning rate is fixed throughout. Experiments are conducted using an NVIDIA Tesla V100 GPU.
IV-B Training Strategy
@@
Crucially, the asynchronous setting is ill-posed for the baseline models as mixing subjects introduces cross-label contamination. We report their asynchronous results (see Table III) only for completeness—to highlight this limitation and to compare against our design, which is expressly built for asynchronous inputs.
IV-C Validation Strategy
Cross Validation
@@ IV-C Validation Strategy
Baselines
@@
We evaluate on both single modality and multimodal fusion baselines. All models inherit similar design of encoders, backbone, and training hyper-parameters as our pipeline to isolates fusion effects from all other factors under both synchronous and asynchronous settings. Single-Modality Baselines: They use the exact architecture from our main pipeline, but only one encoder is active at a time. These serve as upper bounds and isolate modality–specific signal. Conventional Methods: They serve as naive fusion baselines for our approach. (1) Early/Late Fusion: Early Fusion concatenates low-level features from encoders and feeds the joint sequence to the shared backbone; Late Fusion processes each modality independently through the backbone, then concatenates the resulting high-level vectors for prediction. (2) Shared-Latent Fusion: Each modality is linearly projected into a common latent width and fused by element-wise addition before the backbone. (3) Cross-Attention Fusion: A lightweight, symmetric cross-attention block lets each modality attend to the other; the fused sequence is then passed to the backbone.
@@
SOTA Variants: These are implemented to benchmark our method against newer designs. (1) FOCAL [49]: This method factorizes each modality into shared and private latents, enforces orthogonality between them, and applies a contrastive objective to align shared parts across modalities, preserving complementary cues while reducing single-stream shortcuts. (2) TACA [50]: It compresses sequences into a small set of learned tokens per modality and performs cross-attention among tokens (plus a time-shared module) for efficient, salient fusion that tolerates time-misaligned streams. (3) DEEPAV [51]: This method use lightweight per-modality streams followed by late cross-modal interaction and an agreement/consensus head, emphasizing decision-level fusion and robust fallback when one modality is weak or missing.
@@
Dataset Approach w/ class rebalancing w/o class rebalancing Mod.1 Mod.2 Mod.3 Avg Acc Mod.1 Mod.2 Mod.3 Avg Acc FOG Dataset Skeleton IMU — Skeleton IMU — Skeleton Only – – – – – – – – Sensor Only – – – – – – – – Early Fusion – – Late Fusion – – Shared Latent – – Cross Attention – – FOCAL [49] – – TACA [50] – – DEEPAV [51] – – TRIP – – FOG Dataset Skeleton vGRF — Skeleton vGRF — Skeleton Only – – – – – – – – Sensor Only – – – – – – – – Early Fusion – – Late Fusion – – Shared Latent – – Cross Attention – – FOCAL [49] – – TACA [50] – – DEEPAV [51] – – TRIP – – WearGait Dataset Walkway Insole IMU Walkway Insole IMU Walkway Only – – – – – – – – Insole Only – – – – – – – – IMU Only – – – – – – – – Early Fusion Late Fusion Shared Latent Cross Attention FOCAL [49] TACA [50] DEEPAV [51] TRIP
@@
Dataset Approach w/ class rebalancing w/o class rebalancing Acc (%) Acc (%) FOG Skeleton Only – – IMU Only – – Early Fusion Late Fusion Shared Latent Cross Attention FOCAL [49] TACA [50] DEEPAV [51] TRIP WearGait Walkway Only – – Insole Only – – IMU Only – – Early Fusion Late Fusion Shared Latent Cross Attention FOCAL [49] TACA [50] DEEPAV [51] TRIP